OpenStudy (anonymous):

Find an expression in terms of x, tan(arccos(3x)). Thanks.

OpenStudy (anonymous):



OpenStudy (anonymous):

and then the pythagorean thm ?

OpenStudy (anonymous):

need the third side, which your find via pythagoras it is \[\sqrt{1^2-(3x)^2}=\sqrt{1-9x^2}\]

OpenStudy (anonymous):



OpenStudy (anonymous):

then tangent is "opposite over adjacent" and you are done

OpenStudy (anonymous):

it can also be negative right?